OpenStudy (anonymous):

Is (0, 3) a solution to the equation y = x + 3? (1 point) Yes or No

OpenStudy (anonymous):

just sub in your numbers and see... (0,3) x = 0 and y = 3 now just sub it into the equation y = x + 3 3 = 0 + 3 3 = 3 (correct) YES...it is a solution

OpenStudy (anonymous):

Thanks for explaining this!

OpenStudy (anonymous):

no problem :)